OpenStudy (anonymous):

Solve for w. P = 2l + 2w

OpenStudy (johnweldon1993):

\[\large P = 2l + 2w\] We want to isolate 'w' so subtract 2l from both sides of the equation

OpenStudy (anonymous):

\[A.) W=P-l\] \[B.) W= \frac{ P-2l }{ l }\] \[C.)W=\frac{ p-2l }{ 2 }\] \[D.)W=\frac{ 2l-p }{ 2 }\] @johnweldon1993

OpenStudy (johnweldon1993):

Yeah I see the answer up there \[\large P = 2l + 2w\] \[\large P - 2l = 2W\] \[\large W = \frac{P - 2l}{2}\]

OpenStudy (anonymous):

That's what I thought, but I was a bit confused. :)

OpenStudy (johnweldon1993):

Always good to double check :)